Annales Academiæ Scientiarum Fennicæ Mathematica Volumen 31, 2006, 61–70

BANACH SPACES WHOSE BOUNDED SETS ARE BOUNDING IN THE BIDUAL

Humberto Carrión, Pablo Galindo, and Mary Lilian Lourenço

Universidade de São Paulo, Instituto de Matemática e Estatística Dep. de Matemática, Caixa Postal 66281, 05315-970 São Paulo, Brazil; leinad@ime.usp.br Universidad de Valencia, Facultad de Matemáticas Departamento de Análisis Matemático, ES-46100 Burjasot, Valencia, Spain; galindo@uv.es Universidade de São Paulo, Instituto de Matemática e Estatística

Dep. de Matemática, Caixa Postal 66281, 05315-970 São Paulo, Brazil; mllouren@ime.usp.br

Abstract. We discuss necessary conditions for a Banach space to satisfy the property that its bounded sets are bounding in its bidual space. Apart from the classic case of c_{0} , we prove that, among others, the direct sum c_{0}(l_{2}^{n}) is another example of spaces having such property.

A subset B of a complex Banach space X is said to be bounding if every entire function defined on X is bounded on B. Such a set is also a limited set (see below) and, of course, every relatively compact set is bounding. If X is such that every bounding subset is relatively compact, then X is called a Gelfand– Phillips space and there is some literature devoted to the topic. Gelfand–Phillips spaces are characterized also as those whose sequences converging against entire functions are norm convergent [7]. B. Josefson [12] and T. Schlumprecht [15] found simultaneously and independently examples of complex Banach spaces containing limited non-bounding (hence non-relatively compact) sets.

By H(X) we denote the space of entire functions defined on X, that is, functions that are Fréchet differentiable at every point of X. H_{b}(X) is the subspace of all f ∈ H(X) such that f is bounded on bounded sets; these are the so-called entire functions of bounded type.

Bounding sets are related as well to the extension of holomorphic functions to a larger space in the following way: Let E ⊂ F be complex Banach spaces. Every bounded subset of E is bounding in F if, and only if, every f ∈ H(E) having a holomorphic continuation to F is of bounded type. In the particular instance of F = E^{∗∗} it is known that every f ∈ H_{b}(E) has a holomorphic extension to E^{∗∗}. Therefore, in this case the only holomorphic functions which may be extended to the bidual are those of bounded type if, and only if, every bounded set in E is bounding in E^{∗∗}.

2000 Mathematics Subject Classification: Primary 46G20.

The second author was supported by Projects AE-2003-0392 (Universidad de Valencia) and BFM 2003-07540 (DGI. Spain) and the third author in part by CAPES, Brazil, Research Grant AEX0287/03-2.

As far as we know there is essentially one known case (E = c_{0} ) where the above mentioned (equivalent) conditions hold. This was proved among other results by B. Josefson in his 1978 paper [11]. Clearly, no reflexive infinite-dimensional Banach space satisfies those conditions and neither the conditions hold in general for quasi-reflexive Banach spaces, i.e., spaces of finite codimension in its bidual, the conditions may not hold: Consider, for instance, the Q-reflexive (see [1]) and quasi-reflexive Tsirelson–James space T_{J}^{∗}. Its bidual space is separable and hence there the bounding sets are relatively compact (see [5, 4.26]). So the bounded sets in T_{J}^{∗} are not bounding in (T_{J}^{∗})^{∗∗}.

Our aim is to study this situation and to provide further examples of Banach spaces whose bounded sets are bounding in the bidual. So we consider the class ℱ of Banach spaces

ℱ = {F : every bounded set B ⊂ F is a bounding set in F^{∗∗}}.

We show that such a class ℱ is stable under products and quotients, but not under closed subspaces; however the closed subspaces of c_{0} do belong to the class. In the second part of the article we prove that c_{0}(l_{2}^{n}) the predual of C. Stegall’s example of a Schur space whose dual lacks the Dunford–Pettis property [16] is a Banach space whose bounded sets are bounding in the bidual. Our argument does not rely on Josefson’s paper and so provides another proof of the fact that c_{0} is in the class we are dealing with.

1. Generalities

A subset B of a complex Banach space X is bounding if, and only if, lim_{m} ∥P_{m}∥_{B}^{1/m} = 0 for all entire functions f = ∑_{m} P_{m}, where ∑_{m} P_{m} represents the Taylor series of f ∈ H(X) at the origin ([6, Lemma 4.50]). We will use this characterization in Section 2.

A subset A of a Banach space is called limited if weak* null sequences in the dual space converge uniformly on A. The unit ball of an infinite-dimensional Banach space is not a limited set by the Josefson–Nissenzwieg theorem. An operator T: X → Y is said to be limited if it maps the unit ball of X onto a limited set in Y . It turns out that T is limited if, and only if, T^{∗} is weak*-to-norm sequentially continuous. As we mentioned, each bounding set is a limited one, therefore if E ∈ ℱ, then the canonical inclusion E → E^{∗∗} is a limited operator.

Recall that a Banach space X has the property (P) of Pełczyński if for every K ⊂ X^{∗} that is not weakly relatively compact, there exists a weakly Cauchy series ∑ x_{n} in X such that inf_{n} ∥x_{n}∥_{K} > 0. Every C(K) space has the (P) property (see [17, III.D.33]).

We refer to [5] for infinite-dimensional holomorphy background.

Proposition 1.1. If the canonical inclusion E → E^{∗∗} is a limited operator, then E^{∗} is a Schur space. Conversely, if E^{∗} is a Schur space and (i) E^{∗∗} is a Grothendieck space or (ii) E has the (P) property, then the canonical embedding of E into E^{∗∗} is limited.

Proof. If the canonical inclusion E → E^{∗∗} is a limited operator, then the natural projection π_{E^{∗}}: E^{∗∗∗} → E^{∗} is weak*-to-norm sequentially continuous. Therefore the identity on E^{∗} as a composition of the canonical embedding in E^{∗∗∗} and the natural projection is weak-to-norm sequentially continuous, i.e., E^{∗} is a Schur space. For the converse statement (i), note that π_{E^{∗}}: E^{∗∗∗} → E^{∗} is weak-weak sequentially continuous, hence weak*-to-norm sequentially continuous by the assumptions on E^{∗} and E^{∗∗}. And for the (ii) case, π_{E^{∗}}: E^{∗∗∗} → E^{∗} is weak*-to-weak sequentially continuous by [10, Proposition 3.6, p. 131].

Note. If the canonical inclusion E → E^{∗∗} is a limited operator, then H_{b}(E) = H_{wu}(E). This is so because then E^{∗} is a Schur space, hence E has the Dunford–Pettis property and does not contain copies of l_{1} .

In the sequel, we aim at studying whether important classes of Banach spaces belong to the class ℱ.

Firstly, we observe that for no (isomorphic) infinite-dimensional dual space X^{∗}, the canonical inclusion X^{∗} → X^{∗∗∗} can be limited, since otherwise by composing it with the restriction to X operator, ϱ: X^{∗∗∗} → X^{∗}, it would turn out that the identity on X^{∗} would be a limited operator. Therefore, no dual space belongs to ℱ.

When does a C(K)-space belong to ℱ? The above proposition shows that in this setting of Banach spaces of continuous functions the condition of E^{∗} being a Schur space is equivalent to the bounded sets in E being bounding in the bidual: Indeed, if C(K)^{∗} is a Schur space, then C(K) does not contain a copy of l_{1} , hence by [13] its bidual is isomorphic to l_{∞}(I) for some set I. But then C(K)^{∗∗} is a Grothendieck space, so the bounded sets in C(K) are limited in C(K)^{∗∗} and furthermore bounding there because of [11, Theorem 1].

Since c_{0} ∈ ℱ and it is an M-ideal in l^{∞} one wonders whether there are any connections between ℱ and the class of Banach spaces E which are an M-ideal in E^{∗∗}. We are to see that there is no a clear connection.

On one hand, for the Lorentz sequence space d(w, 1) ([10, p. 103]), its predual E = d_{∗}(w, 1) is known to be an M-ideal in its bidual ([10, III 1.4]). However, the unit basis of d(w, 1) is weakly null, so d(w, 1) is not a Schur space. Hence, it turns out that the canonical inclusion E → E^{∗∗} is not a limited operator. Thus, not every bounded set in E is bounding in E^{∗∗}. Another (simpler) example is the following: According to [10, III Examples 1.4(f) and (g)], for 1 < p < ∞, the space K(l_{p}) of compact linear operators of l_{p}, is an M-ideal in its bidual. However K(l_{p})^{∗} is not a Schur space since K(l_{p}) does not have the Dunford–Pettis property since it contains a complemented copy of the reflexive space l_{p}.

On the other hand, the condition of E being an M-ideal in E^{∗∗} is not necessary for their bounded sets being bounding in E^{∗∗}. The Bourgain–Delbaen Banach space Y —introduced in [2]—is shown there to satisfy Y^{∗} ≈ l_{1} and not to contain any copy of c_{0} . Thus Y^{∗∗} ≈ l_{∞} is a Grothendieck space and Proposition 1.1 applies. Further, again by Josefson’s result the limited sets in Y^{∗∗} are bounding sets, so it turns out that the bounded sets in Y are bounding in Y^{∗∗}. However, Y is not an M-ideal in its bidual, since if it were, then it would be isomorphic to c_{0} as a consequence of [10, III Theorem 3.11]. We also mention in passing that Y fails the (P) property.

Corollary 1.2. If the Banach space E is an M-ideal in its bidual, then the canonical inclusion E → E^{∗∗} is a limited operator if, and only if, E^{∗} is a Schur space.

Proof. Since any Banach space which is an M-ideal in its bidual has the (P) property ([10, III Theorem 3.4]), it suffices to apply Proposition 1.1.

Is the class ℱ stable by closed subspaces? The Bourgain–Delbaen space Y also shows that ℱ is not stable by closed subspaces since, as Haydon [9] proved, Y contains a subspace isomorphic to some l_{p}, which clearly does not belong to ℱ.

Recall ([10, III Theorem 1.6]) that every subspace M of K(l_{p}) is also an M-ideal in its bidual and those subspaces which embed into c_{0} satisfy that M^{∗} are Schur spaces ([14, p. 415]). Moreover, every closed subspace E of c_{0} is also an M-ideal in its bidual, has the Dunford–Pettis property and does not contain any copy of l_{1} , hence E has the (P) property and E^{∗} is a Schur space. Therefore, the canonical inclusion E → E^{∗∗} is a limited operator. So the arising question is: Does every closed subspace of c_{0} have its bounded sets bounding in the bidual? Actually this question has a positive answer which reduces to Josefson’s result that c_{0} ∈ ℱ.

Proposition 1.3. The bounded sets in every closed subspace E of c_{0} are bounding in E^{∗∗}.

Proof. To begin with, note that since E has the DP property and does not contain any copy of l_{1} , H_{b}(E) = H_{wu}(E), the space of entire functions on E which are weakly uniformly continuous on bounded sets. Moreover E^{∗} is a separable space, hence by Corollary 12 in [3], H_{wu}(E) = H_{wsc}(E), the space of weakly sequentially continuous entire functions on E.

Let f be entire on E with extension, f̄, to E^{∗∗}. If f ∉ H_{b}(E), f is not weakly sequentially continuous, thus there is a sequence (x_{n}) ⊂ E weakly convergent to a ∈ E such that (f(x_{n})) does not converge to f(a). By considering the function g(x) = f(x + a) we may assume that a = 0. Clearly, (x_{n}) cannot be a norm null sequence, hence by the Bessaga–Pełczyński selection principle (see [4, Chapter V, p. 46]) there is a subsequence of (x_{n}) which is a basic sequence equivalent to a block basic sequence taken with respect to the unit vector basis of c_{0} , and further such subsequence spans a Banach subspace X of E which is complemented in c_{0} along a projection p (see [8, Proposition 249]). There is no loss of generality assuming that the sequence (x_{n}) is such a subsequence. Now f̄|_{X^{∗∗}} is an extension to X^{∗∗} of the function f|_{X}, and therefore, f̄|_{X^{∗∗}} ∘ p^{∗∗} is an extension of f|_{X} ∘ p to l_{∞}. Further, since c_{0} ∈ ℱ, it follows that f|_{X} ∘ p belongs to H_{b}(c_{0}), so it has to be a weakly sequentially continuous function and this is prevented by the choice of (x_{n}) ⊂ X.

We end this section with another positive result.

Proposition 1.4. ℱ is stable by products and quotients.

Proof. (a) Let E, F ∈ ℱ. Let f ∈ H(E^{∗∗}×F^{∗∗}) and consider a bounded set in E×F which we may suppose to be A×B with A ⊂ E and B ⊂ F both bounded sets. We check that the collection {f(x, ·)}_{x∈A} ⊂ H(F^{∗∗}) is τ_{0} bounded: Indeed, for any compact subset K of F^{∗∗}, the collection {f( · , y)}_{y∈K} ⊂ (H(E^{∗∗}), τ_{0}) is bounded, hence is τ_{δ} bounded ([5, 2.44, 2.46]). In addition, since A is bounding in E^{∗∗}, ∥f∥_{A} = sup{|f(x)| : x ∈ A} defines a τ_{δ} continuous seminorm in H(E^{∗∗}) by [5, 4.18], so we have

sup_{x∈A} sup_{y∈K} |f(x, y)| = sup_{y∈K} sup_{x∈A} |f(x, y)| = sup_{y∈K} ∥f( · , y)∥_{A} < ∞,

as we wanted. Now, since B is bounding in F^{∗∗}, we have that {f(x, ·)}_{x∈A} is bounded for the ∥ · ∥_{B} seminorm, hence {|f(x, y)|}_{x∈A,y∈B} is bounded and so A × B is shown to be bounding in E^{∗∗} × F^{∗∗}.

(b) If F is a quotient of a Banach space E ∈ ℱ, then F belongs to ℱ as well: Let q: E → F be the quotient mapping and f ∈ H(F) with a holomorphic extension g to F^{∗∗}. Then g ∘ q^{tt} is a holomorphic function in E^{∗∗} which extends f ∘ q. Therefore, f ∘ q ∈ H_{b}(E) and since any bounded subset A ⊂ F is contained in q(B) for some bounded set B, it follows that f is bounded in A.

2. The example c_{0}(l_{2}^{n})

In this section we consider X = c_{0}(l_{2}^{n}). X^{∗} enjoys the Schur property, yet its dual fails the Dunford–Pettis property, and it was the first example of a space with the Dunford–Pettis property whose dual lacks it, a fact discovered by C. Stegall [16]. Thus in view of Proposition 1.1 we found it suitable to explore whether X belongs to ℱ .

We now set some notation. Consider a partition of N by defining for each n ∈ N the interval

I_{n} = [\frac{n(n − 1)}{2} + 1, \frac{n(n + 1)}{2}] ∩ N.

Stegall’s example is

X^{∗} := {(x_{i}) ∈ l^{∞} : ∑_{n=1}^{∞}(∑_{i∈I_{n}} |x_{i}|^{2})^{1/2} < ∞}.

The canonical predual of X^{∗} is the space

X := {(x_{i}) ∈ l^{∞} : lim_{k→∞} sup_{n≥k}(∑_{i∈I_{n}} |x_{i}|^{2})^{1/2} = 0}.

Observe that X coincides with the closed linear span in X^{∗∗} of the unit vectors {e_{j} = (δ_{nj})_{n∈N}, j ∈ N}.

The dual of X^{∗} can be represented by

X^{∗∗} = {(x_{i}) ∈ l^{∞} : sup_{n∈N}(∑_{i∈I_{n}} |x_{i}|^{2})^{1/2} < ∞}.

Let S be a subset of N. A sequence of complex numbers x = (x_{i}) is said to have support in S if x_{i} = 0 for each i ∉ S. If, moreover, x belongs to a sequence space E, then we write x ∈ E_{S}.

B(E) denotes the closed unit ball of E whatever the normed space E be. For f: E → C we denote ∥f∥_{S} = sup_{x∈B(E_{S})} |f(x)|.

Let (p_{i})_{i∈N} be a strictly increasing sequence of natural numbers. For each i ∈ N, let α_{p_{i}+1}, …, α_{p_{i+1}} be scalars at least one of which is nonzero, and let u_{i} = ∑_{j=p_{i}+1}^{p_{i+1}} α_{j}e_{j} be a block sequence taken from (e_{i})_{i∈N} . The sequence (u_{i})_{i∈N} is called totally disjoint with respect to (I_{n})_{n∈N} whenever p_{i} ∈ I_{n_{i}}, p_{i+1} ∈ I_{n′_{i}} and n_{i} ≠ n′_{i}.

We remark that a block sequence (u_{i})_{i∈N} may be written as u_{i} = ∑_{j∈U_{i}} α_{j}e_{j}, where U_{i} = ⋃_{j∈F_{i}} I_{j} and F_{i} is a finite subset of N. Further, if p_{i} ∈ I_{s}, p_{i+1} ∈ I_{r} such that I_{r} ∩ I_{s} = ∅ we can choose the subsets F_{i} such that max F_{i} < min F_{i+1} and in this case (u_{i})_{i∈N} is a totally disjoint sequence.

In the sequel we consider a sequence (U_{n})_{n∈N} of subsets of N such that for each n

(i) U_{n} = ⋃_{j∈F_{n}} I_{j}, where F_{n} is a finite subset of N and (ii) max F_{n} < min F_{n+1} .

Condition (ii) implies U_{n} ∩ U_{m} = ∅ for each m ≠ n.

It is not difficult to show that if the sequence (u_{i})_{i∈N} , u_{i} = ∑_{j∈U_{i}} α_{j}e_{j}, is totally disjoint and sup_{i∈N} ∥u_{i}∥ ≤ C for some positive constant C, then the partial sums of the series ∑ θ_{i}u_{i} are bounded by C · sup_{i} |θ_{i}| and so the formal series ∑ θ_{i}u_{i} defines an element in X^{∗∗}.

To prove the next theorem we will make use of the following result which extends a similar one from Dineen ([6, p. 299]) valid for l_{∞}.

Lemma 2.1. Let Q be a continuous polynomial in X^{∗∗} such that Q(0) = 0. Let N′ ⊂ N be an infinite set. Then for each ε > 0 there is an infinite subset S ⊂ N′ such that

∥Q∥_{∪_{n∈S}U_{n}} < ε.

Proof. First we consider a homogeneous polynomial P. Suppose that there is an ε > 0 such that for each infinite subset S ⊂ N′ we have

∥P ∥_{∪_{n∈S}U_{n}} > ε.

Let (S_{i})_{i} be a disjoint partition of N′ into an infinite number of infinite sets. Then

∥P ∥_{∪_{n∈S_{i}} U_{n}} > ε for all i ∈ N.

Thus, given i ∈ N there is an x_{i} in X∪_{n∈S_{i}} U_{n}^{∗∗} with ∥x_{i}∥ ≤ 1 such that |P (x_{i})| > ε. Now, by Lemma 1.9 of [6] we have that

sup_{|θ_{i}|=1}|P (∑_{i=1}^{l} θ_{i}x_{i})|^{2} ≥ ∑_{i=1}^{l} |P (x_{i})|^{2} > lε^{2}.

Since the sequence (x_{i}) has pairwise disjoint support, it follows that for |θ_{i}| = 1, i = 1, …, l and any l ∈ N, the combination ∑_{n=1}^{l} θ_{i}x_{i} ∈ B(X^{∗∗}). So, for all l ∈ N we have

∥P ∥^{2} ≥ sup_{|θ_{i}|=1}|P (∑_{i=1}^{l} θ_{i}x_{i})|^{2} ≥ ∑_{i=1}^{l} |P (x_{i})|^{2} > lε^{2}.

This leads to a contradiction when we let l → ∞.

Finally, let Q be an arbitrary continuous polynomial with Q(0) = 0. Then there are homogeneous continuous polynomials P_{j} such that Q = ∑_{j=0}^{m} P_{j} and, by the above, we may find inductively infinite subsets S_{j} ⊂ S_{j−1} ⊂ N′, j = 1, …, m such that ∥P_{j}∥_{∪_{n∈S_{j}} U_{n}} < ε/m. Consequently, ∥Q∥_{∪_{n∈Sm} U_{n}} < ε.

Theorem 2.2. Let (u_{i})_{i∈N} be a totally disjoint sequence in B(X^{∗∗}) given by u_{i} = ∑_{j∈U_{i}} α_{j}e_{j} for each i. Then A = {u_{i} : i ∈ N} is a bounding set in X^{∗∗}.

Proof. Suppose the set A is not bounding in X^{∗∗}. Then there exists an entire function f = ∑_{m} P_{m} ∈ H(X^{∗∗}) such that f is not bounded in A. By Lemma 4.50 in [6] this implies the existence of δ > 0 and strictly increasing subsequences (m_{n}) ⊂ N, (γ_{n}) ⊂ N such that

|P_{γ_{n}}(u_{m_{n}})|^{1/γ_{n}} > δ for all n ∈ N.

We will show that this inequality cannot hold by using an inductive argument.

Given the polynomial y ∈ X^{∗∗} ↦ P_{γ_{1}}(u_{m_{1}} +y)−P_{γ_{1}}(u_{m_{1}}), Lemma 2.1 shows that there is an infinite subset S_{1} ⊂ S_{0} := {m_{n} : n ∈ N} such that

|P_{γ_{1}}(u_{m_{1}} + y) − P_{γ_{1}}(u_{m_{1}})| ≤ (\frac{δ}{γ_{1}!})^{γ_{1}} for all y ∈ B(X∪_{n∈S_{1}} U_{n}^{∗∗}).

Since S_{1} is an infinite subset of S_{0} there is m_{j} ∈ S_{1} with j > 1. Without loss of generality we may suppose that j = 2. Thus u_{m_{2}} ∈ B(X∪_{n∈S_{1}} U_{n}^{∗∗}). Put θ_{1} = 1. Now, by Lemma 1.9(b) in [6], there is θ_{2} ∈ C with |θ_{2}| = 1 such that

|P_{γ_{2}}(θ_{1}u_{m_{1}} + θ_{2}u_{m_{2}})|^{2} ≥ |P_{γ_{2}}(θ_{1}u_{m_{1}})|^{2} + |P_{γ_{2}}(u_{m_{2}})|^{2} ≥ δ^{2γ_{2}}.

The next step is to consider the polynomial y ∈ X^{∗∗} ↦ P_{γ_{2}}(θ_{1}u_{m_{1}}+θ_{2}u_{m_{2}}+y) −P_{γ_{2}}(θ_{1}u_{m_{1}} + θ_{2}u_{m_{2}}). Again Lemma 2.1 shows that there is an infinite subset S_{2} ⊂ S_{1} , such that

|P_{γ_{2}}(θ_{1}u_{m_{1}} + θ_{2}u_{m_{2}} + y) − P_{γ_{2}}(θ_{1}u_{m_{1}} + θ_{2}u_{m_{2}})| ≤ (\frac{δ}{γ_{2}!})^{γ_{2}}

for all y ∈ B(X∪_{n∈S_{2}} U_{n}^{∗∗}). Hence for all y ∈ B(X∪_{n∈S_{2}} U_{n}^{∗∗}),

\begin{aligned}|P_{γ_{2}}(θ_{1}u_{m_{1}} + θ_{2}u_{m_{2}} + y)| & ≥ |P_{γ_{2}}(θ_{1}u_{m_{1}} + θ_{2}u_{m_{2}})| − (\frac{δ}{γ_{2}!})^{γ_{2}} \\ & ≥ δ^{γ_{2}} − (\frac{δ}{γ_{2}!})^{γ_{2}} ≥ (\frac{δ}{2})^{γ_{2}}. \end{aligned}

Arguing by induction we obtain a decreasing sequence (S_{j}) of infinite subsets of N, and a sequence (θ)_{i}, θ_{1} = 1, in the unit sphere of C such that

|P_{γ_{j}}(∑_{k=1}^{j} θ_{k}u_{m_{k}} + y)| ≥ (\frac{δ}{2})^{γ_{j}} for all y ∈ B(X∪_{n∈S_{j}} U_{n}^{∗∗}).

We set a = (a_{l}) ∈ X^{∗∗} defined by a_{l} = θ_{k}α_{l}e_{l} if l ∈ U_{m_{k}} for some k ∈ N, and a_{l} = 0 otherwise. That is, a is the formal series ∑_{k=1}^{∞} θ_{k}u_{m_{k}}.

Since the sets U_{k} are disjoint, the vectors u_{k} have disjoint supports, so a ∈ B(X^{∗∗}) and also a−∑_{k=1}^{j} θ_{k}u_{m_{k}} ∈ B(X∪_{n∈S_{j}} U_{n}^{∗∗}). Then from the latter inequality we conclude

|P_{γ_{j}}(a)|^{1/γ_{j}} = |P_{γ_{j}} (∑_{k=1}^{j} θ_{k}u_{m_{k}} + (a − ∑_{k=1}^{j} θ_{k}u_{m_{k}}))|^{1/γ_{j}} ≥ \frac{δ}{2}.

Since {a} is a compact set in X^{∗∗}, it follows that lim_{j→∞} |P_{γ_{j}}(a)|^{1/γ_{j}} = 0. A contradiction.

To prove the next theorem we need the following lemma.

Lemma 2.3. Let (u_{i})_{i∈N} be a totally disjoint bounded sequence given by u_{i} = ∑_{j=p_{i}+1}^{p_{i+1}} α_{j}e_{j} for each i ∈ N. Then there exists a projection Q: X → \overline{span[u_{i}]}.

Proof. Let (u′_{i})_{i∈N} be a sequence in X^{∗} such that ∥u′_{i}∥ = 1, u′_{i}u_{i} = 1 for each i ∈ N and u′_{i}(e_{j}) = 0 for j ∉ {p_{i} + 1, …, p_{i+1}}.

If we define Q(x) = ∑_{i=1}^{∞} u′_{i}(x)u_{i} for each x = (x_{j}) ∈ X, since u′_{i}(x) = u′_{i}(∑_{j=p_{i}+1}^{p_{i+1}} x_{j}e_{j}), we have that

|u′_{i}(x)| ≤ ∥ ∑_{j=p_{i}+1}^{p_{i+1}} x_{j}e_{j}∥ ≤ 2Λ∥∑_{i=1}^{∞} ∑_{j=p_{i}+1}^{p_{i+1}} x_{j}e_{j}∥ ≤ 2Λ∥x∥,

where Λ is the basis constant for (e_{n})_{n∈N} . Then

∥∑_{i=1}^{∞} u′_{i}(x)u_{i}∥ = ∥∑_{i=1}^{∞} u′_{i}( ∑_{j=p_{i}+1}^{p_{i+1}} x_{j}e_{j})u_{i}∥ ≤ sup_{i}|u′_{i}( ∑_{j=p_{i}+1}^{p_{i+1}} x_{j}e_{j})| ≤ 2Λ∥x∥.

So Q is a continuous projection.

Theorem 2.4. Every bounded subset of X is bounding in X^{∗∗}.

Proof. Notice that X^{∗} is separable. So, by a result of [3], it is sufficient to prove that every weakly compact subset of X is bounding in X^{∗∗}. So by Eberlein’s theorem, it suffices to show that every weakly convergent sequence (x_{i}) ⊂ X is a bounding set in X^{∗∗}. We may assume that (x_{i}) is weakly null since bounding sets are also bounding after translation.

Suppose that (x_{i}) is not a bounding subset in X^{∗∗}. Then, there is an entire function f = ∑_{n∈N} P_{n} ∈ H(X^{∗∗}) which is unbounded on the set {x_{i} : i ∈ N}. Thus, by Lemma 4.50 in [6] there exist a subsequence of (x_{i}) (which we are going to denote in the same way), a subsequence (P_{γ_{i}}), and δ > 0 such that

|P_{γ_{i}}(x_{i})|^{1/γ_{i}} > δ.

Since (x_{i}) is a weakly null non null sequence, we find by the Bessaga– Pełczynski selection principle, a subsequence (x_{k_{i}}) equivalent to a basic block sequence taken from (e_{i}). That is, there is a strictly increasing sequence (p_{i}) ⊂ N such that (x_{k_{i}}) ≈ (u_{i}), where u_{i} = ∑_{j=p_{i}+1}^{p_{i+1}} α_{j}e_{j} for i ∈ N. Since (p_{i}) is strictly increasing we may assume, passing to subsequences if necessary, that (u_{i}) is totally disjoint, and henceforth a bounding set in X^{∗∗}.

Since (x_{k_{i}}) is equivalent to (u_{i}), there exists an isomorphism T from [u_{i}, i ∈ N] onto [x_{k_{i}}, i ∈ N] such that T (u_{i}) = x_{k_{i}} for each i ∈ N. Let T^{∗∗}: [u_{i}, i ∈ N]^{∗∗} ⟶ [x_{k_{i}}, i ∈ N]^{∗∗} be the double transpose of T and let Q^{∗∗}: X^{∗∗} ⟶ [u_{i}, i ∈ N]^{∗∗} be the double transpose of the projection defined in Lemma 2.3. Since f ∈ H(X^{∗∗}) we have that f ∘ T^{∗∗} ∘ Q^{∗∗} ∈ H(X^{∗∗}) and f ∘ T^{∗∗} ∘ Q^{∗∗}|_{[u_{i}]} coincides with f ∘ T ∘ Q.

Finally, since {u_{i} : i ∈ N} is a bounding subset in X^{∗∗} by Theorem 2.2, we get that lim_{i} |P_{γ_{k_{i}}} ∘ T^{∗∗} ∘ Q^{∗∗}(u_{i})|^{1/γ_{k_{i}}} = 0 for each i ∈ N, but lim_{i} |P_{γ_{k_{i}}} ∘ T^{∗∗} ∘ Q^{∗∗}(u_{i})|^{1/γ_{k_{i}}} = lim_{i} |P_{γ_{k_{i}}} ∘ T ∘ Q(u_{i})|^{1/γ_{k_{i}}} = lim_{i} |P_{γ_{k_{i}}} (x_{k_{i}})|^{1/γ_{k_{i}}} . So, lim_{i} |P_{γ_{k_{i}}} (x_{k_{i}})|^{1/γ_{k_{i}}} = 0. This contradicts our assumption and completes the proof.

70 H. Carrión, P. Galindo, and M.L. Lourenço

Acknowledgments. Part of this paper was done while the third-named author was visiting the Departamento de Análisis Matemático at University of Valencia. She thanks this department for its hospitality and specially Professor Pablo Galindo.

The question we have dealt with arose some years ago after a cup of coffee with Professor Seán Dineen to whom we are warmly grateful.

We thank the referee for his valuable suggestions which led to improvements of the paper.

References

[1] Aron, R., and S. Dineen: Q-Reflexive Banach spaces. - Rocky Mountain J. Math. 27:4, 1997, 1009–1025.

[2] Bourgain, J., and F. Delbaen: A class of special ℒ_{∞} spaces. - Acta Math. 145, 1981, 155–176.

[3] Carrión, H.: Entire functions on Banach spaces with a separable dual. - J. Funct. Anal. 189, 2002, 496–514.

[4] Diestel, J.: Sequences and Series in Banach Spaces. - Springer, New York, 1984.

[5] Dineen, S.: Complex Analysis on Locally Convex Spaces. - North-Holland Math. Stud. 57, 1981.

[6] Dineen, S.: Complex Analysis on Infinite Dimensional Spaces. - Springer, 1999.

[7] Galindo, P., L. A. Moraes, and J. Mujica: Weak holomorphic convergence and bounding sets in Banach spaces. - Proc. Roy. Irish Acad. 98A:2, 1998, 153–157.

[8] Habala, P., P. Hajek, and V. Zizler: Introduction to Banach Spaces. - Charles University, Prague, 1997.

[9] Haydon, R.: Subspaces of the Bourgain–Delbaen space. - Studia Math. 139, 2000, 275– 293.

[10] Harmand, P., D. Werner, and W. Werner: M-ideals in Banach Spaces and Banach Algebras. - Lecture Notes in Math. 1547, Springer, 1993.

[11] Josefson, B.: Bounding subsets of l_{∞}(A). - J. Math. Pures Appl. 57, 1978, 397–421.

[12] Josefson, B.: A Banach space containing non-trivial limited sets but no non-trivial bounding set. - Israel J. Math. 71, 1990, 321–327.

[13] Pełczyński, A., and Z. Semadeni: Spaces of continuous functions (III). - Studia Math. 18, 1959, 211–222.

[14] Saksman, E., and H.-O. Tylli: Structure of subspaces of the compact operators having the Dunford–Pettis property. - Math. Z. 232, 1999, 411–425.

[15] Schlumprecht, T.: A limited set that is not bounding. - Proc. Roy. Irish Acad. 90A, 1990, 125–129.

[15] Stegall, C.: Duals of certain spaces with the Dunford–Pettis property. - Notices Amer. Math. Soc. 19, 1972, 799.

[16] Wojtasczyk, P.: Banach Spaces for Analysts. - Cambridge Studies in Advanced Mathematics, Cambridge Univ. Press, 1991.

Received 14 October 2004